OpenStudy (anonymous):

Find all real solutions (x+2)^2=(x-4)^2

OpenStudy (raden):

first, expand for LHS and RHS be : (x+2)^2=(x-4)^2 x^2 + 4x + 4 = x^2 - 8x + 16

OpenStudy (raden):

move side all terms RHS to LHS : x^2 + 4x + 4 - x^2 + 8x - 16 = 0 just combine the similar terms : 12x - 12 = 0 solve for x

OpenStudy (anonymous):

thank you I forgot my special products and thought it was x^2+4= x^2 - 16

OpenStudy (raden):

note : (a+b) ^ 2 = a^2 + 2ab + b^2 and (a-b) ^ 2 = a^2 - 2ab + b^2 welcome ..

OpenStudy (anonymous):

:D

OpenStudy (raden):

memorize it :D

OpenStudy (anonymous):

o_- I shall